OpenStudy (duckymomo):

Susan's math teacher assigned x homework problems on Monday. On Tuesday, she assigned 14 more problems. Over the two days, a total of 22 homework problems were assigned. Which equation could be used to find x, the number of problems assigned on Monday?

OpenStudy (hoblos):

x+14 = 22

OpenStudy (anonymous):

22-14=8

OpenStudy (hoblos):

so x=22-14=8

OpenStudy (duckymomo):

x + 14 = 22 22 + x = 14 x - 14 = 22 14x = 22

OpenStudy (duckymomo):

thats a b c and d

OpenStudy (hoblos):

a